As a financial analyst, you are tasked with finding the price per share of ABC, a telecom company.

PartA - For this part, assume that the Gordon growth model for stock valuation holds. The typical company in the telecom industry has a ratio of next period's dividend to current price D1/P0 = 2%, pays out 40% of earnings as dividends, and generates a return on equity of 15%. Your analysis suggests that the discount rate of the typical company in the telecom industry is appropriate to value ABC's stock. What is this discount rate?

PartB - For this part of the analysis, you assume that the differential dividend growth approach can be used to correctly value ABC's stock. ABC just paid $1.60 per share in dividends. You estimate that dividends will grow at 15% for the next three years until year 3, then at 7% perpetually. Find the current price per share of ABC.

PartC - The expected earnings of ABC for next year is $2.20 per share. What is the net present value per share of ABC's growth opportunities?

PartD - You recompute the price of ABC assuming instead that dividends will always grow at a constant rate g and obtain the same price you got in part (b). What dividend growth rate did you assume?
Answer #1

1.
required return=dividend yield+RoE*(1-payout ratio)=2%+15%*(1-40%)=11.000%

2.
Share price=Present value of dividends

=1.6*1.15/1.11+1.6*(1.15/1.11)^2+1.6*(1.15/1.11)^3+1.6*(1.15/1.11)^3*1.07/(11%-7%)=52.75010145

3.
PVGO=Share Price-Expected Earnings/required return

=52.75010145-2.20/11%=32.75010145

4.
Share Price=Expected Dividend/(required return-growth rate)

52.75010145=1.60*(1+g)/(11%-g)
=>g=7.73229751670702%